OpenStudy (anonymous):

Two angles of a quadrilateral each measure 70°. A third angle of the quadrilateral measures 115°. What is the measure of the fourth angle of the quadrilateral? A. 40° B. 65° C. 105° D. 115°

OpenStudy (anonymous):

i think the answer is d

OpenStudy (anonymous):

plz help

OpenStudy (nikato):

No Remember all 4 angles of a quadrilateral add up to 360

OpenStudy (anonymous):

its c

OpenStudy (mokeira):

2(70)+115=255 360-255=105

OpenStudy (anonymous):

thank you all you 4 your help

OpenStudy (mokeira):

The answer is C

OpenStudy (anonymous):

Because equi has total angle = 360 360-70-70-115=105

OpenStudy (anonymous):

Two angles of a triangle measure 58° and 104°. What is the measure of the third angle of the triangle?

OpenStudy (midhun.madhu1987):